# The Feynman paradox in a spherical axion insulator

### v1

@@ III.2 Angular momentum of the electromagnetic field
The integral corresponding to the potentials outside the sphere has to be split into two contributions, one for the region $a<R<d$ and another one for the region $R>a$, due to the expansions (15) and (16). Similarly to the planar case [Nogueira-van_den_Brink_PhysRevResearch.4.013074], it is easy to show that the second integral in Eq. (III.2) corresponding to the interior of the TI vanishes and thus we actually have,
@@
where the prime denotes a derivative with respect to $\cos\Theta$. The above expression can be simplified further by using the orthogonality of Legendre polynomials along with the identities,
@@
with $\mathcal{B}_{n}=\beta_{n}/q\alpha$, and $B_{n}=b_{n}/q$. We will use the latter constants to leading order in $\alpha$, since there is already a prefactor $\alpha^{2}$ in Eq. (53). We have,
@@
Note that the angular momentum (via the function $\Upsilon$) is only a function of $d/a$, although this is not immediately obvious from the series of Eq. (54). The angular momentum as a function of $d/a$ and $\theta=\pi$ is shown in Fig. 2. It is interesting to observe that $L_{z}$ can be sizable via e.g., $N$. Unlike Eq. (III.1), the dependence on $N$ is not linear, but quadratic. Thus, in a typical STM tip one usually has $N\sim 10^{2}$ [STM], implying that actually measuring the angular momentum should not represent a major technical challenge. For $d\gg a$, when the charge is very far away from the TI, Eq. (54) turns into,
@@
Figure 3: The $d$-dependence of the ratio of the electromagnetic angular momentum $L_{z}^{\infty}$ when $d\gg a$, to the angular momentum of the Hall current $L_{\text{Mech},z}$, with different values of $N$. The gray dashed line crosses points, where the ratio is equal to unity.
@@
In Fig. 3 we show the $d$-dependence of the ratio of the electromagnetic angular momentum in Eq. (57) to the one in Eq. (III.1). Though the analytical expressions in the above mentioned equations are not identical, we can determine a regime where the ratio is equal to unity numerically. In first order of $\alpha$ this ratio depends only on $N$ and $d$, and does not on either $\epsilon$ or $a$. Varying the former ones, we get the family of points corresponding to the transfer of angular momentum to the surface Hall current.
Finally, it is worth to mention that Eq. (57) includes only odd powers of the fine-structure constant, while Eq. (III.1) – only even powers of $\alpha$.
We may wonder what happens in the so called ‘ìnfinite sphere” limit, where the radius $a$ of the sphere is much larger than $d$, as discussed in Ref. [Lindell-AJP]. This case would correspond to the one of a semi-infinite TI, whose angular momentum due to the electromagnetic field was calculated in Ref. [Nogueira-van_den_Brink_PhysRevResearch.4.013074]. The result turns out to be independent of the distance of the charge $q$ to the surface, in a situation reminiscent of the angular momentum of a dyon (a dipole consisting of an electric charge and a magnetic monopole). Indeed, the angular momentum of a dyon is independent of its size. However, for a slab of finite thickness, the dependence on the distance of the charge to the surface arises in the angular momentum. An example of such a calculation is found in Appendix A.
@@ Acknowledgements.
We acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), through SFB 1143 project A5 and the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter-ct.qmat (EXC 2147, Project Id No. 390858490), as well as support by Bundesministerium für Forschung, Technologie und Raumfahrt (BMFTR) funding through project No. 01DK240008 (GU-QuMat).
@@ Appendix A Infinite slab of thickness $L$
For semi-infinite magnetoelectric discussed in detail in Ref. [Nogueira-van_den_Brink_PhysRevResearch.4.013074], the $z$-component of electromagnetic angular momentum turns out to not depend on the distance $d$ from the magnetoelectric surface to the point-charge placed above,
@@
and $\Phi_{0}=hc/e$ is the flux quantum, $\kappa=(1/2)(\alpha\theta/\pi)^{2}$. Now, to investigate the origin of the $d$-dependence of the angular momentum, we turn to an intermediate example of an infinite slab of thickness $L$ (Fig. 4), since such geometry can be viewed as stereographic projection of the sphere onto the plane. In this case, the north and south poles become the upper and lower parts of the slab, respectively.
Figure 4: Infinite magnetoelectric slab of thickness $L$ discussed in Appendix A.
The system of equations in part IV of the Ref. [Nogueira-van_den_Brink_PhysRevResearch.4.013074] is extended by one region z<Lz<-L,

### v2

@@ III.2 Angular momentum of the electromagnetic field
The integral corresponding to the potentials outside the sphere has to be split into two contributions, one for the region $a<R<d$ and another one for the region $R>a$, due to the expansions (15) and (16). Similarly to the planar case [13], it is easy to show that the second integral in Eq. (III.2) corresponding to the interior of the TI vanishes and thus we actually have
@@
where the prime denotes a derivative with respect to $\cos\Theta$. The above expression can be simplified further by using the orthogonality of Legendre polynomials along with the identities
@@
with $\mathcal{B}_{n}=\beta_{n}/q\alpha$ and $B_{n}=b_{n}/q$. We will use the latter constants to leading order in $\alpha$, since there is already a prefactor $\alpha^{2}$ in Eq. (53). We have,
@@
Note that the angular momentum (via the function $\Upsilon$) is only a function of $d/a$, although this is not immediately obvious from the series of Eq. (54). The angular momentum as a function of $d/a$ and $\theta=\pi$ is shown in Fig. 2. It is interesting to observe that $L_{z}$ can be sizable via e.g., $N$. Unlike Eq. (III.1), the dependence on $N$ is not linear, but quadratic. Thus, in a typical STM tip one usually has $N\sim 10^{2}$ [21], implying that actually measuring the angular momentum should not represent a major technical challenge. For $d\gg a$, when the charge is very far away from the TI, Eq. (54) turns into
@@
Figure 3: The $d$ dependence of the ratio of the electromagnetic angular momentum $L_{z}^{\infty}$ in Eq. (57) when $d\gg a$, to the angular momentum of the Hall current $L_{\text{Mech},z}$, with different values of $N$. The gray dashed line crosses points, where the ratio is equal to unity.
@@
In Fig. 3 we show the $d$ dependence of the ratio of the electromagnetic angular momentum in Eq. (57) to the one in Eq. (III.1). Though the analytical expressions in the above mentioned equations are not identical, we can determine a regime where the ratio is equal to unity numerically. In first order of $\alpha$ this ratio depends only on $N$ and $d$, and does not on either $\epsilon$ or $a$. Varying the former ones, we get the family of points corresponding to the transfer of angular momentum to the surface Hall current.
Finally, it is worth to mention that Eq. (57) includes only odd powers of the fine-structure constant, while Eq. (III.1) only even powers of $\alpha$.
We may wonder what happens in the so called ‘ìnfinite sphere” limit, where the radius $a$ of the sphere is much larger than $d$, as discussed in Ref. [9]. This case would correspond to the one of a semi-infinite TI, whose angular momentum due to the electromagnetic field was calculated in Ref. [13]. The result turns out to be independent of the distance of the charge $q$ to the surface, in a situation reminiscent of the angular momentum of a dyon (a dipole consisting of an electric charge and a magnetic monopole). Indeed, the angular momentum of a dyon is independent of its size. However, for a slab of finite thickness, the dependence on the distance of the charge to the surface arises in the angular momentum. An example of such a calculation is found in the Appendix.
@@ Acknowledgements.
We acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), through SFB 1143 Project A5 and the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter-ct.qmat (EXC 2147, Project ID No. 390858490), as well as support by Bundesministerium für Forschung, Technologie und Raumfahrt (BMFTR) funding through Project No. 01DK24008 (GU-QuMat).
@@ Appendix A Infinite slab of thickness $L$
For semi-infinite magnetoelectric discussed in detail in Ref. [13], the $z$ component of electromagnetic angular momentum turns out to not depend on the distance $d$ from the magnetoelectric surface to the point-charge placed above,
@@
and $\Phi_{0}=hc/e$ is the flux quantum, $\kappa=(1/2)(\alpha\theta/\pi)^{2}$. Now, to investigate the origin of the $d$ dependence of the angular momentum, we turn to an intermediate example of an infinite slab of thickness $L$ (Fig. 4), since such geometry can be viewed as stereographic projection of the sphere onto the plane. In this case, the north and south poles become the upper and lower parts of the slab, respectively.
Figure 4: Infinite magnetoelectric slab of thickness $L$ discussed in Appendix.
The system of equations in part IV of Ref. [13] is extended by one region z<Lz<-L,